# AI-Driven Skin Cancer Diagnosis: Grad-CAM and Expert Annotations for Enhanced Interpretability

### v1

@@ 3.2 Clinical-inspired XAI: BCC diagnosis with additional label information
This section analyzes the performance of the AI tool for BCC detection in conjunction with the labels provided to explain this classification. Table 3 presents metrics that summarize this performance. The metrics are averaged over all folds. This table has three parts. The first part shows the performance of the AI tool in the binary classification. The second part shows its performance in detecting BCC dermoscopic patterns. Finally, the third part represents the accuracy of the labels that provide the clinical explanation.
@@
Overall, the BCC/non-BCC diagnostic performance is high, around $0.9$ for all metrics. However, the BCC pattern detection performance has to be analysed with a deeper insight. Minority classes tend to attain low recall because the AI tool trained with unbalanced databases tends to favor majority classes. As shown in Sect. 2.1, SW, MG and ML are underrepresented classes. Strategies such as data augmentation and advanced sampling, a one-vs-all strategy combined with stratified k-fold cross-validation helped to achieve a more balanced classification across patterns, thereby improving overall model performance. However, the metrics achieved should not be analyzed in the same way as BCC/non-BCC performance. They should only be evaluated to the extent that they provide a correct explanation for the binary classification. It is not relevant if the AI tool misses a specific BCC pattern, but if it misses any BCC pattern, as clinicians diagnose skin lesions in the same way. This further evaluation is summarized in the third part of Table 3. As shown in this table, 73 percent of non-BCC lesions without any BCC pattern, 95 percent of non-BCC lesions with PN, and 99 percent of BCC lesions with some BCC pattern are correctly labeled as such.
@@
Table 3: Evaluation using binary and multilabel classification metrics, fine-tuned binary classifier, and physician-guided analysis.
Recall Specificity Precision Accuracy
@@
Pigment Network 0.94 0.96 0.97 0.95
Ulceration 0.81 0.75 0.52 0.77
OvidNests 0.65 0.84 0.53 0.84
Multiglobules 0.61 0.81 0.32 0.80
Maple Leaf-like 0.50 0.82 0.34 0.77
Spoke Wheel 0.60 0.87 0.37 0.84
Arborizing Telangiectasia 0.89 0.76 0.61 0.80
Clinical-inspired XAI
@@
3.3 Clinical-inspired Visual XAI
This section aims to quantify the accuracy of the AI tool in focusing on the correct part of the lesion, specifically the BCC dermoscopic patterns identified by clinicians. To this end, BCC pattern areas delineated by dermatologists will be compared with model activated areas. This will provide a quantitative measure of the model’s agreement with human diagnostic criteria and demonstrate its ability to accurately identify critical features of BCC lesions.
[…]
Fig. 4 illustrates this analysis. Fig. 4(a) shows the original BCC lesion. Fig. 4(b) shows the Grad-CAM map. Fig. 4(c) shows the dermatologist’s segmentation overlaid on the Grad-CAM map. Fig. 4(d) shows an example of the two conditional probability density functions. The orange curve represents $P\left(z\left(x,y\right)\mid\text{Bg}\right)$, and the blue curve represents $P\left(z\left(x,y\right)\mid\text{Fg}\right)$. The orange curve is centered near 0, indicating low activation outside the mask, while the blue curve shows significant Grad-CAM information within the clinical segmentation, indicating that the model extracts features from the same region as the specialist.
Table 4: Statistics derived from estimation of conditional probability density functions of GradCAM within and outside the region of clinical interest.
Prediction Intersection Mean Fg Mean Bg Std Fg Std Bg
@@ 3.3 Clinical-inspired Visual XAI
(a) Original image
(b) GradCAM
(c) Dermatologist segmentation
(d) $P\left(z\left(x,y\right)\mid\text{Fg}\right)$ and $P\left(z\left(x,y\right)\mid\text{Bg}\right)$
Figure 4: Clinically-inspired visual interpretation of a correct prediction.
Table 4 summarizes the information extracted from these probability density function. Specifically, mean, standard deviation of $z\left(x,y\right)$ for $(x,y)\in Fg$ and $(x,y)\in\text{Bg}$ respectively, and the intersection area between $P\left(z\left(x,y\right)\mid\text{Fg}\right)$ and $P\left(z\left(x,y\right)\mid\text{Bg}\right)$ are shown. This table shows that correctly predicted samples have a larger mean standard deviation than incorrectly predicted samples. In addition, the intersection area is larger in these cases. These facts prove that the model is not able to pay attention to the areas of clinical interest in the incorrect predictions.
4 Discussion and conclusion
In this paper, an AI tool for BCC diagnosis that provides a clinical explanation has been developed. It achieves two main achievements that make it very clinically significant. First, this tool will contribute to changing the protocol of teledermatology, reducing the waiting time for diagnosis and intervention in an area of 60 geographically dispersed primary health centers. Second, its clinically inspired double explanation increases its usefulness.

### v2

@@ 3.2 Clinically-inspired XAI: BCC diagnosis with additional label information
This section analyses the performance of the AI tool for BCC detection in conjunction with the labels provided to explain this classification. Table 3 presents metrics that summarise this performance. The metrics are averaged over all folds. This table has three parts. The first part shows the performance of the AI tool in binary classification. The second part shows its performance in detecting BCC dermoscopic patterns. Finally, the third part represents the accuracy of the labels that provide the clinical explanation.
@@
In general, the diagnostic performance of BCC / non-BCC is high, around $0.9$ for all metrics. However, the BCC pattern detection performance needs to be analysed with a deeper insight. Minority classes tend to achieve low recall because the AI tool trained with unbalanced databases tends to favour majority classes. As shown in Sect. 2.1, SW, MG and ML are underrepresented classes. Strategies such as data augmentation and advanced sampling, a one-vs-all strategy combined with stratified k-fold cross-validation helped to achieve a more balanced classification across patterns, thereby improving overall model performance. However, the metrics achieved should not be analysed in the same way as BCC/non-BCC performance. They should only be evaluated to the extent that they provide a correct explanation for binary classification. For the specialist, a tool that perfectly classifies all BCC patterns within a lesion is not particularly useful, as identifying even a single pattern can already indicate a positive BCC diagnosis. Therefore, the third section of Table 3 provides information relevant to the clinician, highlighting whether there is a BCC pattern detected by the specialist is present. Furthermore, this information is further supported by Table 4, which, as will be discussed later, shows whether the classification of each pattern is being extracted from the same region as identified by the dermatologist. This additional evaluation is summarised in the third section of Table 3, titled “Clinically-inspired XAI”. The table demonstrates that 73% of non-BCC lesions without any BCC patterns, 95% of non-BCC lesions with PN, and 99% of BCC lesions with at least one BCC pattern are correctly labelled as such.
@@
Table 3: Evaluation using binary, multilabel classification metrics and physician-guided analysis.
Recall ($\sigma^{2}$) Specificity ($\sigma^{2}$) Precision ($\sigma^{2}$) Accuracy ($\sigma^{2}$)
@@
Pigment Network 0.94 ($4.19\cdot 10^{-4}$) 0.96 ($1.07\cdot 10^{-4}$) 0.97 ($2.47\cdot 10^{-4}$) 0.95 ($4.50\cdot 10^{-6}$)
Ulceration 0.81 ($4.14\cdot 10^{-3}$) 0.75 ($8.94\cdot 10^{-4}$) 0.52 ($8.94\cdot 10^{-4}$) 0.77 ($4.84\cdot 10^{-4}$)
OvidNests 0.65 ($2.48\cdot 10^{-3}$) 0.84 ($1.52\cdot 10^{-4}$) 0.53 ($3.78\cdot 10^{-5}$) 0.84 ($1.20\cdot 10^{-5}$)
Multiglobules 0.61 ($6.16\cdot 10^{-3}$) 0.81 ($2.05\cdot 10^{-3}$) 0.32 ($2.15\cdot 10^{-3}$) 0.80 ($1.10\cdot 10^{-3}$)
Maple Leaf-like 0.50 ($8.76\cdot 10^{-3}$) 0.82 ($9.43\cdot 10^{-5}$) 0.34 ($1.79\cdot 10^{-3}$) 0.77 ($2.47\cdot 10^{-4}$)
Spoke Wheel 0.60 ($2.63\cdot 10^{-3}$) 0.87 ($8.09\cdot 10^{-4}$) 0.37 ($1.15\cdot 10^{-3}$) 0.84 ($3.93\cdot 10^{-4}$)
Arborizing Telangiectasia 0.89 ($1.08\cdot 10^{-3}$) 0.76 ($4.57\cdot 10^{-4}$) 0.61 ($4.69\cdot 10^{-4}$) 0.80 ($2.83\cdot 10^{-4}$)
Clinically-inspired XAI
@@
3.3 Clinically-inspired Visual XAI
This section aims to quantify the accuracy of the AI tool in focusing on the correct part of the lesion, specifically the BCC dermoscopic patterns identified by clinicians. To this end, the areas of the BCC pattern delineated by dermatologists will be compared with the activated areas of the model. This will provide a quantitative measure of the model’s agreement with human diagnostic criteria and demonstrate its ability to accurately identify critical features of BCC lesions.
To quantify the accuracy of the model activation areas with respect to the areas of clinical interest, the conditional probability density functions of the normalised Grad-CAM values within and outside the area segmented by dermatologist were estimated. Let $z(x,y)$ the Grad-CAM value at position $(x,y)$. Denote Fg the area segmented by the dermatologist and Bg the background. $P\left(z\left(x,y\right)\mid\text{Fg}\right)$ is the probability density function of the Grad-CAM values for pixels $(x,y)\in\text{Fg}$ and $P\left(z\left(x,y\right)\mid\text{Bg}\right)$ is the probability density function of Grad-CAM values for pixels $(x,y)\in\text{Bg}$.
Fig. 4 illustrates this analysis. Fig. 4(a) shows the original BCC lesion. Fig. 4(b) shows the Grad-CAM map. Fig. 4(c) shows the dermatologist’s segmentation overlaid on the Grad-CAM map. Fig. 4(d) shows an example of the two conditional probability density functions. The orange curve represents $P\left(z\left(x,y\right)\mid\text{Bg}\right)$, and the blue curve represents $P\left(z\left(x,y\right)\mid\text{Fg}\right)$. The orange curve is centred near 0, indicating low activation outside the mask, while the blue curve shows significant Grad-CAM information within the clinical segmentation, indicating that the model extracts features from the same region as the specialist.
Table 4 summarises the information extracted from these probability density functions. Specifically, the mean, standard deviation of $z\left(x,y\right)$ for $(x,y)\in Fg$ and $(x,y)\in\text{Bg}$ respectively, and the intersection area between $P\left(z\left(x,y\right)\mid\text{Fg}\right)$ and $P\left(z\left(x,y\right)\mid\text{Bg}\right)$ are shown. This table highlights that correctly predicted samples exhibit a higher average $P\left(z\left(x,y\right)\mid\text{Fg}\right)$ value compared to misclassified samples. In addition, they display a lower average $P\left(z\left(x,y\right)\mid\text{Bg}\right)$ value, which suggests that the model is focusing on regions akin to those a dermatologist would examine when diagnosing the lesion.
The DICE and Jaccard coefficients are commonly used to evaluate accuracy. However, in this study, these metrics were not applicable due to the significant difference in detail between manual segmentations by specialists and the more generalized regions highlighted by Grad-CAM (Fig. 3). Activation maps tend to cover larger areas than the precise segmentations, making these coefficients less representative for assessing model performance.
(a) Original image and the detailed BCC pattern segmentation by a specialist in red.
(b) GradCAM heatmap showcasing areas of highest model activation.
(c) Integration of Grad-CAM heatmap and specialist segmentation.
(d) Analysis of Grad-CAM heatmap and expert segmentation. The blue density curve indicates that a significant portion of the Grad-CAM is accurately segmented by the expert.
Figure 4: Demonstration of various interpretative visualizations used in skin lesion analysis, combining expert segmentation and Grad-CAM heatmaps to aid in understanding the model’s decision-making process. This example is from a correct prediction sample.
Table 4: Statistics derived from estimation of conditional probability density functions of Grad-CAM within and outside the region of clinical interest.
Prediction Intersection Mean $P\left(z\left(x,y\right)\mid\text{Fg}\right)$ Mean $P\left(z\left(x,y\right)\mid\text{Bg}\right)$ Std $P\left(z\left(x,y\right)\mid\text{Fg}\right)$ Std $P\left(z\left(x,y\right)\mid\text{Bg}\right)$
@@ 3.3 Clinically-inspired Visual XAI
4 Discussion
The aim of this work was to develop a tool that is truly useful to dermatologists by incorporating clinically relevant information, such as BCC patterns and their segmentation. The tool provides not only a BCC/non-BCC classification but also a deeper level of interpretability through detailed pattern analysis. However, further refinement is needed to ensure that the tool can deliver not only a binary diagnosis of BCC/non-BCC but also a comprehensive identification of specific patterns with accuracy percentages for each. This level of detail aims to make the information truly valuable for teledermatology applications.
The two essential criteria for the clinical diagnosis of BCC by dermatologists are: (1) the presence or absence of any BCC pattern, and (2) the Pigment Network as a negative criterion. The Clinically-inspired XAI tool correctly detects 99% of cases where any of the six BCC patterns are present in the lesion. Additionally, we aimed to verify that the tool adheres closely to clinical criteria, which is why the double explanation approach was implemented, as illustrated in the results.
As demonstrated in Table 3, the clinical reliability of the diagnosis was confirmed to be robust for specialists. Furthermore, Table 4 indicates that the feature extraction for each dermoscopic pattern was consistently performed from clinically relevant regions, aligning the model’s attention with the areas identified by dermatologists. This alignment validates the model’s ability to not only distinguish between BCC and non-BCC lesions but also to provide precise pattern-based explanations that improve interpretability and clinical trust.
In the literature, there are other studies that have developed tools aimed at addressing similar challenges. Nevertheless, the tool presented in this work uniquely combines BCC/non-BCC detection with an explainability component for BCC patterns, which has been clinically validated as shown in Table 4. This additional layer of clinical interpretability provides a distinctive advantage that existing tools lack, significantly enhancing its value in a medical context.
The implementation of this tool will allow the development of a new primary care protocol in Andalusia, streamlined the workflow between primary care and dermatologists. Its deployment would be highly efficient, as the entire system is built using the lightweight and efficient MobileNet model, which will facilitate this implementation. In addition, its strong clinical reliability provides a solid foundation for adoption by specialists, ensuring both accuracy and ease of use in clinical settings.
5 Acknowledgements
This work was supported by the Andalusian Regional Government (PROYEXCEL_00889) and is part of the project of the aid PID2021-127871OB-I00, funded by MCIN/AEI/10.13039/501100011033 and ERDF/European Union (NextGenerationEU/PRTR).
Conflict of Interest Statement
The authors declare that they have no conflict of interest regarding this study.